OpenStudy (anonymous):

The graphs of r = cos theta and r = 2 cos 2 theta intersect in how many points?

OpenStudy (anonymous):

I know one will be a rose and one will be a circle but I don't know how to calculate intersection

OpenStudy (campbell_st):

find the points of intersection.... by solving alternatively... graph both curves 2cos(2x) takes pi for 1 period... while cos(x) takes 2pi for a period.... so they intersect twice over (0, pi) or 4 times over (0, 2pi)

OpenStudy (campbell_st):

This is an equation reducible to a quadratic 2cos(2x) = 2(cos^2x - sin^2(x) using sin^2 x = 1 - cos^2(x) then 2cos(2x) = 2(2cos^2(x) - 1) the point of intersection is found by equating both sides 2(2cos^2(x) -1) = cos(x) 4cos^2(x) - cos(x) - 2 = 0 solve the quadratic using the general quadratic formula cos(x) =\[(1 \pm \sqrt{32})/8\]

OpenStudy (campbell_st):

oops should be \[\cos(x) = (1 \pm \sqrt{33})/8\]

OpenStudy (campbell_st):

I'll leave you to determine the domain of x